LOFAR Newsletters April 2016

Distributed on 26 April 2016

 

Array status:

----------------------

-       38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

-       12 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609, PL610, PL611, PL612.

-       3 new international stations have been built in Poland (PL610, PL611, and PL612). Final checks are in progress to successfully include them in all ILT observations (see below).

-       In 2016, a new LOFAR station will be built on the grounds of Birr Castle, located centrally in Ireland.

-       Oscillating tiles at a few stations were detected during BF observations and were deactivated.

-       CS031 had an issue with strange polarization behaviour in mode 1 (LBA_OUTER). The reason was traced to a wrong calibration table being applied by mistake. This is now corrected.

-       The overview of non-operational antenna elements for LBA and HBA is available here. Maintenance will start within a few weeks.

-       Station calibration: data have been collected for LBA outer and sparse even. The automatic reduction has finished and manual corrections are now to be performed before the tables can be installed. HBA mode 5 has also been collected for core and remote stations. COBALT delays need to be estimated with the new tables before these can be put in place. PL610 data in LBA show the presence of RFI sources that prevent successful station calibration. Investigations are in progress.

 

Observing System Status:

------------------------------------ 

-       The observing system was mostly stable during the period March - April 2016. However, failures of COBALT nodes, and the instability of processing and LTA (ingest) systems caused some delays in the exploitation of the Cycle 5 observing programme and in digesting data out of CEP2.

-       The LTA database went down on 15 March. It has been restored after a few days. The issue affected ingest, retrieval and, query on the database.

-       Successful observations including the new Polish stations have been performed and first maps have been generated (see here). However, the inclusion of the stations in any HBA DUAL observing setup along with the rest of the array highlighted frequent drops in sensitivity for a random number of Dutch stations. A tiger team is currently active to identify the cause of the problem and solve it.

 

Software Development Status (J. Annyas):

------------------------------------------

-       Software version 2.16.2 was rolled out on 13 April. This added support for CEP4.

-       CEP4: Validation is underway. Manually started observations and pipelines are successful. Current focus is on finalization of the server configuration and integration with the LOFAR operational software.

-       PSNC (Poznan) has been commissioned and has become an active LTA storage facility.

-       Beta version of Ingest pipeline is successfully tested with EOR transfers.

 

CITT Status (E. Orru' & T. J. Dijkema)

--------------------------------------------------------------------

- The CITT2 is currently focusing on stress testing the Factor pipeline, the pipeline for HBA imaging using the facet calibration technique.

  • Busy Week 24 took place at Astron  (22-25 February). About 10 fields passed the initial stages of this pipeline. 
  • The pipeline is now able to process interleaved observations.
  • The functionality of multi-scale selfcalibration has been added to FACTOR in order to make more gradual the process of producing 5 arcsec resolution images.


- In DPPP the option to solve for multiple channels and to correct for rotation angle has been implemented and is currently being tested.

- Software able to interface the Image Domain Gridder to various imaging tools is under development.

 

Observing Programmes:

--------------------------------------------------------------------

-       30 regular Cycle 6 proposals were received at the deadline of 9 March 2016. The review process is currently ongoing and will conclude with the PC meeting on 28+29 April, at ASTRON.

-       Cycle 5 observing programme: 86% complete. Schedule available here.

-       Cycle 4 observing programme: 95% complete. The rest is being observed with second priority during Cycle 5.

 

Documentation:

--------------------------------------------------------------------

-       A collection of LOFAR slides that can be used when preparing LOFAR-related presentations is available here.

-       Recent LOFAR science highlights are collected here and in ppt form here.

-       An updated list of LOFAR papers is available here.

 

CEP news:

----------------

-       CEP4

  • See software development status section.

 

-       CEP2:

  • Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.
  • During the past two months, a few locus nodes experienced issues and were temporarily taken out of production.
  • Locus096 taken out of production and used for spare parts.

 

-       CEP3:

  • Cluster info and schedule available here.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.

 

Meetings:

--------------------------------------------------------------------

-       'The Third LOFAR Users meeting',  ‘'The 2016 LOFAR Community Science Workshop, and the ‘ILT Technical Operations Meeting' took place on 4-7 April 2016 in Zandvoort aan Zee, the Netherlands. The presentations given at the meetings are available here.

-       The 4th LOFAR Data Processing School will be held at ASTRON on 5-9 September 2016. The School website is here. The registration closes on 1 June.

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

Note: the following events are now marked on an online calendar that is available here.

-       Next LSM's : 11/05, 25/05, 08/06, 22/06

(note: all presentations given at the LSM are available here).

-       Next Stop days: 10/05, 07/06

-       Next software roll outs: 23/05

-       4th LOFAR Data Processing School: 5-9/09

-       Next LOFAR bulletin: June 2016.

Design: Kuenst.    Development: Dripl.    © 2021 ASTRON